The sum of two numbers is 30. Their difference is 6. What are the two numbers?
Answer by checkley71(8403)

X+Y=30 & X-Y=6 THEN X=6+Y OR (6+Y)+Y=30 OR 6+2Y=30 OR 2Y=24 OR Y=12 THEN
X+12=30 OR X=30-12 OR X=18
PROOF 18+12=30 & 18-12=6